6. You are called in as a financial analyst to appraise the bonds of Olsen’s Clothing Stores. The $1,000 par value bonds have a quoted annual interest rate of 13 percent, which is paid semiannually. The yield to maturity on the bonds is 12 percent annual interest. There are 20 years to maturity. Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

a. Compute the price of the bonds based on semiannual analysis. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Bond price

b. With 15 years to maturity, if yield to maturity goes down substantially to 10 percent, what will be the new price of the bonds? (Do not round intermediate calculations. Round your final answer to 2 decimal places.

10. Beasley Ball Bearings paid a $4 dividend last year. The dividend is expected to grow at a constant rate of 4 percent over the next four years. The required rate of return is 16 percent (this will also serve as the discount rate in this problem). Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods.

g. If current EPS were equal to $5.33 and the P/E ratio is 1.2 times higher than the industry average of 6, what would the stock price be? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Stock price

h. By what dollar amount is the stock price in part g different from the stock price in part f? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Amount

i. With regard to the stock price in part f, indicate which direction it would move if:

(1) D1 increases
(2) Ke increases
(3) g increases

1.
Answer #1

As per rules I am answering the first 4 subparts of the question

A; Using financial calculator

Input: FV= 1000;

I/Y=12/2=6

N=20*2=40

PMT=13%*1000/2 = 65

Solve for PV as -1,075.23

Price = $1,075.23

B: Using financial calculator

Input: FV= 1000;

I/Y=10/2=5

N=15*2=30

PMT=13%*1000/2 = 65

Solve for PV as -1,230.59

Price = $1,230.59

10 Stock price= D1/(1+k)+ D2/(1+k)^2 + (D3) /(1+k)^3+ D4/(1+k)^4

= 4*(1+4%)^1/1.16^1+4*(1+4%)^2/1.16^2+4*(1+4%)^3/1.16^3+4*(1+4%)^4/1.16^4

=12.27

G: Stock price= EPS*PE Ratio

= 5.33*1.2*6

=38.38